OpenStudy (anonymous):

a tugboat goes 180 miles upstream in 15hrs. the return trip downstream takes 5hrs. Find the speed of the tugboat w/o a current and the speed of the current.

OpenStudy (lgbasallote):

upstream speed is given as u = b - c downstream speed is given as d = b + c where u = upstream speed b = speed of boat c = speed of curent d = downstream speed do you follow so far?

OpenStudy (anonymous):

i follow. what do I do next ?

OpenStudy (lgbasallote):

now.. can you identify what the upstream speed is?

OpenStudy (anonymous):

yes, the upstream speed is 12mph right ?

OpenStudy (lgbasallote):

yes right

OpenStudy (lgbasallote):

how about downstream speed?

OpenStudy (anonymous):

the downstream speed is 36mph right ?

OpenStudy (lgbasallote):

right again.

OpenStudy (lgbasallote):

so you have 12 = b - c an 26 = b + c are you familiar with systems of linear equations?

OpenStudy (anonymous):

sort of

OpenStudy (lgbasallote):

do you know the elimination method?

OpenStudy (anonymous):

yes. I think the answer is 24. Is that correct ?

OpenStudy (lgbasallote):

yes! speed of boat = 24

OpenStudy (lgbasallote):

now find speed of current

OpenStudy (anonymous):

how ?

OpenStudy (lgbasallote):

rememeber: d = b + c you have the value of downstream speed and the speed of the boat

OpenStudy (anonymous):

ok, so the speed of the boat is 24 which makes the speed of the current 12 right ?

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so there's your answer b = 24 mph c = 12 mph

OpenStudy (anonymous):

ok, thanks

OpenStudy (lgbasallote):

welcome